# COMPASS: Computational Mapping of Patient-Therapist Alliance Strategies with Language Modeling

### v1

@@ Abstract
The therapeutic working alliance is a critical factor in predicting the success of psychotherapy treatment. Traditionally, working alliance assessment relies on questionnaires completed by both therapists and patients. In this paper, we present COMPASS, a novel framework to directly infer the therapeutic working alliance from the natural language used in psychotherapy sessions. Our approach utilizes advanced large language models to analyze transcripts of psychotherapy sessions and compare them with distributed representations of statements in the working alliance inventory. Analyzing a dataset of over 950 sessions covering diverse psychiatric conditions, we demonstrate the effectiveness of our method in microscopically mapping patient-therapist alignment trajectories and providing interpretability for clinical psychiatry and in identifying emerging patterns related to the condition being treated. By employing various neural topic modeling techniques in combination with generative language prompting, we analyze the topical characteristics of different psychiatric conditions and incorporate temporal modeling to capture the evolution of topics at a turn-level resolution. This combined framework enhances the understanding of therapeutic interactions, enabling timely feedback for therapists regarding conversation quality and providing interpretable insights to improve the effectiveness of psychotherapy.
@@ 1 Introduction
The working alliance, which encompasses various cognitive and emotional aspects of the therapist-patient relationship, is a critical concept in psychotherapy identified as a crucial factor in predicting treatment outcomes [1, 2]. However, current methods for assessing the alliance rely on evaluating entire therapy sessions using point-scale ratings [3]. These methods can be time-consuming and lack scalability for analyzing individual dialogue turns. Psychotherapy was among the first disciplines to embrace Natural Language Processing (NLP), with early applications including chatbots like ELIZA and Parry, which simulated psychotherapists and patients with schizophrenia, respectively [4, 5]. However, none of these attempts have been adopted or even systematically researched.
@@
For evidence-based psychotherapies such as cognitive behavioral therapy (CBT) [6], quantifying best practices on a large scale has proven challenging. While CBT has demonstrated efficacy for various mental health conditions, including anxiety and depression, capturing the nuances of effective therapeutic techniques in real-world settings remains complex [7]. Traditional assessment methods fall short in providing the necessary granularity and scalability to capture the subtleties of therapist-patient interactions within CBT sessions, which can be important during child and adolescent developmental stages [8].
@@
Following recent advancements in NLP [9, 10, 11, 12, 13], we propose here an approach for quantifying patient-therapist alliance by leveraging large language models (LLMs) to project each dialogue turn onto representations of established working alliance inventories [3, 14, 15]. Our approach enables us to not only estimate the overall degree of alliance but also identify fine-grained patterns and dynamics across shorter and longer time scales, e.g., turn-by-turn in a session and across sessions.
@@
We employ the Counseling and Psychotherapy Transcripts from the Alexander Street dataset [16], which includes transcribed recordings of over 950 therapy sessions between multiple anonymized therapists and patients. We introduce the Working Alliance Transformer (WAT) and the Working Alliance LSTM (WA-LSTM), a Transformer-based and a Long-Short Term Memory-based model, respectively, that leverage the Working Alliance Inventory (WAI) [3, 11], and demonstrate how our method enhances the accuracy of classification models for identifying psychiatric conditions based on therapy transcripts. Previous studies have demonstrated that NLP techniques, including topic modeling, can reveal latent structures within depression-related language collected from platforms like Twitter [17] and improve the detection of Post-Traumatic Stress Disorder [18]; to facilitate interpretable insights, we conduct a systematic investigation of the most prominent topics that are addressed at the turn level (Figure 1).
@@ 2.1 Deep Learning Inference of Working Alliance
The analytic framework for inferring the working alliance from psychotherapy sessions is illustrated in Figure 1. Our approach involves processing the full records of individual patients with multiple clinical conditions or a cohort of patients with the same condition, which can be segmented based on timestamps or topic turns. The original data is presented in pairs of dialogues, and we extract features in three different ways: (1) using the full pairs of dialogues, (2) extracting only the patients’ responses, or (3) extracting only the therapist’s responses. Each feature set has its advantages and disadvantages. The dialogue features contain all the information but can mix together the intents within sentences from both individuals. The patient features provides a more coherent narrative, but it only represent part of the overall story. The therapist features, which can be seen as a type of semantic labeling of the patient’s feelings, can be informative in terms of the diagnostics, but may oversimplify the complexity of the interaction.
The dialogue between the patient and therapist in a session is transcribed into pairs corresponding to the patient’s turn, followed by the therapist’s turn 111This is of course under the assumption that, therapists often provide responses that are broad and generalized, affirming or summarizing the patient’s input. These responses can be thought of as semantic “labels” that can be anticipated based on the “inputs” of the patient’s statements. In reality, the patient-doctor dialogue could as well be initiated by the doctor. For the analytical purposes, we set the default to be patient first, without loss of generalizability to a lag of at most one turn.. The inventories of working alliance questionnaires are also provided in pairs for the patient and the therapist, each comprising 36 statements. We employ sentence or paragraph embeddings to encode both the dialogue turns and the inventories; the embeddings are vectorial representations of text [19] that we then use to compute the similarity between turns and inventory item. This approach yields a 36-dimensional inferred working alliance score for each patient and therapist turn; we will further discuss the specific scales of our inferred working alliance scores in Section 2.3 (see also Appendix Supplementary Material for additional details).
@@ 2.2 Sentence Embeddings
To represent the dialogue turns and working alliance inventories, we employ deep sentence embeddings. In this study, we use two types of sentence embeddings, Doc2Vec and SentenceBERT, which are two popular choices of deep learning-based embedding models of documents or sentences. We used these two models of different neural architectures to demonstrate the model-agnostic feasibility of the microscopic linguistic analytics.
@@ 2.3 Working Alliance Inventories
The Working Alliance Inventory (WAI) is a self-report measurement questionnaire designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [3, 14, 15]. It has been widely used to assess the quality of the working alliance between therapists and patients and has demonstrated good psychometric properties [14, 15]. The modern version of the WAI consists of 36 questions, and participants are asked to rate each item on a 7-point scale (1=never, 7=always) [15]. The inventory aims to measure alliance factors across different types of therapy, establish the relationship between the alliance measure and theoretical constructs underlying the measure, and relate the alliance measure to a unified theory of therapeutic change [23].
@@
The 36 items of the WAI are used to derive three alliance scales: the task scale, the bond scale, and the goal scale. These scales capture the collaborative nature of the patient-therapist relationship, the affective bond between therapist and patient, and the agreement on treatment-related tasks and long-term goals [23]. Each scale score is computed using a weighting matrix that assigns weights to the questionnaire responses based on a key table, resulting in a comprehensive assessment of the working alliance. Additionally, the overall working alliance score is obtained by summing the scores of the three scales [23].
@@ 2.4 Identification of Working Alliance Interaction Patterns
Specifically, we concatenate the 36-dimensional working alliance scores estimated from the current turn, as described above, with the unbiased sentence embedding of the turn. This combined feature vector is then fed into a sequence classifiers, which we term Working Alliance Transformer (WAT) and Working Alliance LSTM (WA-LSTM) (see Suppl. Mat.Supplementary Material). By applying the WAT or the WA-LSTM to the psychotherapy transcripts, we can classify the clinical condition of the sequence based on the working alliance scores and the content of the dialogue turns. This classification model can be applied to the entirety of a session or a segment of the session; its accuracy provides a validation that these patterns can be identified, as we show in the Results section.
@@ 2.5 Psychotherapy Topic Modeling Framework
Using a language model to interpret text data offers several advantages: (1) the model can provide an objective analysis of the data, devoid of personal biases or preconceived notions that human researchers might inadvertently introduce; (2) by relying on the model’s interpretation, researchers can access a more neutral perspective, enhancing the objectivity of their findings; (3) language models can quickly process and analyze large volumes of text, identifying patterns, relationships, and insights that may be challenging for humans to detect manually. This not only saves time but also broadens the scope of interpretability, enabling researchers to explore more nuanced aspects of the data.

### v2

@@ Abstract
The therapeutic working alliance is a critical factor in predicting the success of psychotherapy treatment. Traditionally, working alliance assessment relies on questionnaires completed by both therapists and patients. In this paper, we present COMPASS, a novel framework to directly infer the therapeutic working alliance from the natural language used in psychotherapy sessions. Our approach utilizes advanced large language models (LLMs) to analyze transcripts of psychotherapy sessions and compare them with distributed representations of statements in the working alliance inventory. Analyzing a dataset of over 950 sessions covering diverse psychiatric conditions including anxiety, depression, schizophrenia, and suicidal tendencies, we demonstrate the effectiveness of our method in providing fine-grained mapping of patient-therapist alignment trajectories and offering interpretability for clinical psychiatry and in identifying emerging patterns related to the condition being treated. By employing various deep learning-based topic modeling techniques in combination with prompting generative language models, we analyze the topical characteristics of different psychiatric conditions and their evolution at a turn-level resolution. This combined framework enhances the understanding of therapeutic interactions, enabling timely feedback for therapists regarding the quality of therapeutic relationships and providing interpretable insights to improve the effectiveness of psychotherapy.
@@ 1 Introduction
The working alliance, which encompasses various cognitive and emotional aspects of the therapist-patient relationship, is a critical concept in psychotherapy identified as a crucial factor in predicting treatment outcomes [1, 2]. Research over the past few decades has consistently demonstrated the significant role that a strong working alliance plays in determining the success of various therapeutic modalities, especially in patient engagement and adherence to treatment [3, 4, 5]. However, current methods for assessing the alliance rely on evaluating entire therapy sessions using point-scale ratings [6], with questionnaires and clinical instruments such as the working alliance inventory (WAI). These methods can be time-consuming and lack scalability for analyzing individual dialogue turns. Psychotherapy was among the first disciplines to embrace Natural Language Processing (NLP), with early applications including chatbots like ELIZA and Parry, which simulated psychotherapists and patients with schizophrenia, respectively [7, 8]. However, none of these attempts have been adopted or even systematically researched.
As a current standard in the field, the working alliance inventory (WAI) is a self-report measurement designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [6, 9, 10]. It is typically completed by both patient and therapist at multiple points during therapy to monitor the development and quality of the therapeutic alliance. Common practice involves administering the WAI at the beginning of treatment to establish a baseline, at mid-treatment to track progress, and toward the end to assess changes in the alliance over time. In some settings, it may also be completed periodically throughout therapy, such as after every few sessions or even after each session, to continuously evaluate alliance strength.
The WAI has been widely used to assess the quality of the working alliance between therapists and patients, demonstrating strong psychometric properties [9, 10]. Research indicates that higher WAI scores are associated with better therapy outcomes across various diagnoses, including depression and anxiety disorders. The effect sizes for the correlation between alliance quality and treatment outcomes range from 0.19 to 0.32, suggesting a moderate relationship [11, 12]. The WAI’s reliability is evidenced by a high internal consistency (Cronbach’s alpha around 0.93) and validation across diverse populations and therapeutic settings [11, 13].
Alliance development patterns can vary based on the therapeutic approach. For instance, in cognitive-behavioral therapy (CBT), a strong initial alliance is often crucial for patient engagement and treatment adherence [13]. Studies have also noted differences in alliance strength and its associations with treatment outcomes across diagnostic groups. A study comparing alliance ratings among patients with depression, somatoform disorders, and eating disorders found no significant differences in the strength of alliance ratings or their associations with treatment outcomes across these groups. All three groups reported positive alliances that improved over the course of therapy [14]. In all studied disorder groups, there was a notable incongruence between patient and therapist ratings of the therapeutic alliance. This discrepancy highlights the importance of aligning perceptions to enhance therapeutic processes and outcomes. Clients with severe mental illnesses, such as schizophrenia or personality disorders, often rate the alliance higher than their therapists do [15]. This discrepancy can impact therapy dynamics and outcomes [15]. For instance, clients with substance misuse issues tend to exhibit larger rating discrepancies compared to those with other severe disturbances. Patterns of alliance development can differ based on diagnosis, with some studies identifying stable alliances or linear growth patterns as being correlated with better outcomes [11]. Understanding these patterns can help tailor therapeutic approaches to individual needs.
Our study leverages the WAI as a foundation for analyzing patient-therapist interactions, specifically examining the trajectory of alliance scores over time across different psychiatric conditions. By applying automated analysis methods to infer alliance scores at a granular, turn-by-turn level, we extend the traditional use of WAI. This approach enables us to explore fine-grained dynamics that are typically overlooked in standard WAI applications. Such an analysis can provide actionable insights into how alliance trajectories may vary by diagnosis, offering a deeper understanding of how therapeutic bonds develop within sessions. Ultimately, this could inform tailored therapeutic strategies that align more closely with the unique needs of each patient group, enhancing the utility of WAI in diverse clinical contexts.
@@
For evidence-based psychotherapies such as cognitive behavioral therapy (CBT) [16], quantifying best practices on a large scale has proven challenging. While CBT has demonstrated efficacy for various mental health conditions, including anxiety, depression, and psychosis, capturing the nuances of effective therapeutic techniques in real-world settings remains complex [17]. Traditional assessment methods fall short in providing the necessary granularity and scalability to capture the subtleties of therapist-patient interactions within sessions, which can be important, particularly, during child and adolescent developmental stages [18]. The therapeutic alliance in CBT not only enhances engagement but also facilitates collaborative empiricism, a key principle where both therapist and client actively explore thoughts and behaviors together [19]. By examining alliance patterns within sessions, therapists can make real-time adjustments to interventions, refining techniques to meet individual needs. Hence, understanding this process via alliance can allow us to better quantify the best practice to generalize across populations. This is similarly true in other therapy settings, such as psychodynamic therapy and psychotherapy in general. For instance, in psychodynamic therapy, which often delves into unconscious processes and past experiences, the therapeutic alliance is particularly vital. The quality of this alliance can significantly affect the patient’s willingness to engage with challenging material related to their emotions and relationships. A positive alliance can facilitate the exploration of transference, where feelings about significant others are projected onto the therapist—enabling patients to gain insights into their relational patterns [11].
@@
Despite its utility, the WAI is limited by its scope and intermittent administration, which may miss within-session processes that are central to effective therapy, such as the moment-to-moment shifts in alliance that drive therapeutic progress. Traditional tools like the WAI provide valuable insights but often lack the granularity to capture these intricate dynamics in real-time. To address this gap, we propose here an approach for quantifying patient-therapist alliance by leveraging large language models (LLMs) to project each dialogue turn onto representations of established working alliance inventories [6, 9, 10]. Following recent advancements in NLP [20, 21, 22, 23, 24], our approach enables us to not only estimate the overall degree of alliance but also identify fine-grained patterns and dynamics across shorter and longer time scales, e.g., turn-by-turn in a session and across sessions.
@@
We employ the Counseling and Psychotherapy Transcripts from the Alexander Street dataset [25], which includes transcribed recordings of over 950 therapy sessions between multiple anonymized therapists and patients with clinical diagnoses of anxiety, depression, schizophrenia, or suicidal tendencies. We introduce the Working Alliance Transformer (WAT) and the Working Alliance LSTM (WA-LSTM), a Transformer-based and a Long-Short Term Memory-based model, respectively, that leverage the Working Alliance Inventory (WAI) [6, 22], and demonstrate how our method identifies patterns in therapy transcripts discriminative of the different psychiatric conditions. Previous studies have demonstrated that NLP techniques, including topic modeling, can reveal latent structures within depression-related language collected from platforms like Twitter [26] and improve the detection of Post-Traumatic Stress Disorder [27]; to facilitate interpretable insights, we conduct a systematic investigation of the most prominent topics that are addressed at the turn level (Figure 1).
@@ 2 Methods and Materials
2.1 Psychotherapy Transcript Dataset
We begin by introducing the dataset used in our study. The Alexander Street Counseling and Psychotherapy Transcripts dataset [25] consists of transcribed recordings of over 950 therapy sessions between multiple anonymized therapists and patients. Each session covers one of four psychiatric conditions: anxiety, depression, schizophrenia, and suicidal tendencies. This comprehensive collection includes speech-translated transcripts of the recordings from real therapy sessions, 40,000 pages of client narratives, and 25,000 pages of reference works.
The Alexander Street data includes a large repository of anonymized transcripts of therapy sessions, representing a wide range of approaches including prominently Group Counseling, Cognitive Behavior Therapy, Person-centered Therapy, Brief Relational Therapy, Relational Emotive Behavior Therapy, Psychoanalysis, and Integrative Psychoanalysis, spanning sessions conducted from 1970 through 2012, with client ranging in age from 11 to 80+ yo, an approximate female-male distribution of 70%-30% and less than 10% bisexual/non-binary gender; the majority of therapists/counselors are anonymous (75%+).
This dataset provides a comprehensive source for analyzing the therapeutic process in psychotherapy, where we extracted the following cohort with 498 sessions for anxiety, 377 for depression, 71 for schizophrenia, and 12 for suicidal tendencies. The sessions cover four types of psychiatric conditions: anxiety, depression, schizophrenia, and suicidal. Each dialogue pair consists of a patient response turn followed by a therapist response turn . In total, the dataset contains over 200,000 turns from both patients and therapists, providing a rich source for analyzing the therapeutic process in psychotherapy.
@@ 2.2 Deep Learning Inference of Working Alliance
The analytic framework for inferring the working alliance from psychotherapy sessions is illustrated in Figure 1. Our approach involves processing the full records of individual patients with multiple clinical conditions or a cohort of patients with the same condition, which can be segmented based on timestamps or topic turns. The original data is presented in pairs of dialogues, and we extract features in three different ways: (1) using the full pairs of dialogues, (2) extracting only the patients’ responses, or (3) extracting only the therapists’ responses. Each feature set has its advantages and disadvantages. The dialogue features contain all the information but can mix together the intents within sentences from both individuals.
The dialogue between the patient and therapist in a session is transcribed into pairs corresponding to the patient’s turn, followed by the therapist’s turn 111This is of course under the assumption that, therapists often provide responses that are broad and generalized, affirming or summarizing the patient’s input. These responses can be thought of as semantic “labels” that can be anticipated based on the “inputs” of the patient’s statements. In reality, the patient-doctor dialogue could as well be initiated by the doctor. For the analytical purposes, we set the default to be patient first, without loss of generalizability to a lag of at most one turn.. The inventories of working alliance questionnaires are also provided in pairs for the patient and the therapist, each comprising 36 statements. We employ sentence or paragraph embeddings to encode both the dialogue turns and the inventories; the embeddings are vectorial representations of text [28] that we then use to compute the similarity between turns and inventory item. This approach yields a 36-dimensional inferred working alliance score for each patient and therapist turn; we will further discuss the specific scales of our inferred working alliance scores in Section 2.4 (see Supplementary Materials).
@@ 2.3 Sentence Embeddings
To represent the dialogue turns and working alliance inventories, we employ deep sentence embeddings. In this study, we use two types of sentence embeddings, Doc2Vec and SentenceBERT, which are two popular choices of deep learning-based embedding models of documents or sentences. We used these two models of different neural architectures to demonstrate the model-agnostic feasibility of the fine-grained linguistic analytics.
@@
| Example sentences | Inferred BOND | Inferred TASK | Inferred GOAL |
| 1. “I feel really understood and supported by you. It’s comforting to know I can talk about anything here without being judged.” | 1.41 | -0.76 | -0.66 |
| 2. “I understand that practicing these relaxation techniques as daily tasks will help me manage my anxiety better.” | -1.39 | 0.46 | 0.93 |
| 3. “I’m committed to working towards better communication with my partner, as we discussed in our last session.” | -0.85 | -0.56 | 1.40 |
@@ 2.4 Working Alliance Inventories
The Working Alliance Inventory (WAI) is a self-report measurement questionnaire designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [6, 9, 10], typically completed by patient and therapist at multiple points during therapy to monitor the development and quality of the therapeutic alliance. Common practice involves completing it at the beginning of treatment to establish a baseline, mid-treatment to track progress, and toward the end to assess changes in the alliance over time. In some cases, it may also be completed periodically throughout therapy, such as after every few sessions, or even after each session to continuously evaluate the strength of the alliance.
It has been widely used to assess the quality of the working alliance between therapists and patients and has demonstrated good psychometric properties [9, 10]. Research indicates that higher WAI scores are generally associated with better therapy outcomes across various diagnoses, including depression and anxiety disorders. Effect sizes for the correlation between alliance quality and treatment outcomes range from 0.19 to 0.32, suggesting a moderate relationship [11, 12]. The effectiveness of the therapeutic alliance can vary depending on the type of therapy employed. For example, cognitive-behavioral therapy (CBT) often shows different patterns of alliance development compared to psychodynamic approaches. Studies suggest that in CBT, a strong initial alliance may be crucial for engagement and adherence to treatment [13]. The WAI has demonstrated high internal consistency (Cronbach’s alpha around 0.93) and has been validated across diverse populations and settings [11, 13].
@@
The modern version of the WAI consists of 36 questions, and participants are asked to rate each item on a 7-point scale (1=never, 7=always) [10], as detailed by the full inventory in the supplementary material (Table S1). The working alliance inventories are usually administered with two versions, a client / patient version, and a therapist version. The inventory items are very similar with only a minor change of phrasing. The inventory aims to measure alliance factors across different types of therapy, establish the relationship between the alliance measure and theoretical constructs underlying the measure, and relate the alliance measure to a unified theory of therapeutic change [3]. Both patients and therapists fill out the WAI independently after each session, and responses are blinded to the other party to ensure unbiased assessments.
@@
The 36 items of the WAI are used to derive three alliance scales: Task (the agreement on therapy-related tasks), Bond (the affective bond between therapist and patient), and Goal (the agreement on treatment goals). These scales capture the collaborative nature of the patient-therapist relationship, the affective bond between therapist and patient, and the agreement on treatment-related tasks and long-term goals [3]. Example items include: “The therapist and I agree on what is important for me to work on” (Task), “I feel that the therapist likes me” (Bond), and “We agree on the steps to be taken to improve my situation” (Goal). Each scale score is computed using a weighting matrix that assigns weights to the questionnaire responses to these inventory items based on a key table, resulting in a comprehensive assessment of the working alliance. Additionally, the overall working alliance score is obtained by summing the scores of the three scales [3].
The assessment of WAI is largely restricted to research studies and has not been adopted in the regular practice of therapy, given it is relatively burdensome for both patients and therapist. Even in studies, it is not feasible for the participants to provide analysis of alliance at the turn level, something usually not considered or assessed by third parties. In contrast, we implemented such turn-by-turn analysis, relying on the proven capabilities of large language models to identify nuanced semantic content [32], including the possibility of mapping self-evaluated inventories onto semi-structured interviews and monologues [33].
To illustrate how the process works, here’s an example of three possible utterances occurring in a psychotherapy session (Table 1). We can see the first sentence is clearly discussing the BOND aspect of the therapeutic relationship, where the client feels understood and supported by the therapist. The second sentence highlights the TASK aspect, focusing on the daily tasks that will help the client manage their anxiety as a GOAL. The third sentence emphasizes the GOAL aspect, reflecting the client’s commitment to working towards better communication with their partner.
To quantify these aspects, we first compute their sentence embeddings in a high-dimensional space (300 dimensions in this case, using a Doc2Vec model). For instance, the first sentence’s embedding is computed to be [-0.11, 0.05, 0.08,…], the second sentence become [0.05, 0.02, 0.12,…], and the third [-0.01, 0.10, -0.08,…]. Additionally, we have the working alliance inventories (36 items/questions) in their respective 300-dimension embedding space. For example, the first item "I felt uncomfortable with _." corresponds to [0.01, 0.02, 0.01,…].
We then compute the sentence similarity (e.g., cosine similarity) of the 300-dimension embedding vector of the example sentence with the 300-dimension embedding vector of the inventory item, yielding a 36-dimensional similarity score. For example, the first sentence corresponds to [0.52, 0.49, 0.47, …], with 0.52 being the similarity between the first sentence and the first inventory item. Using the key table corresponding to the working alliance scores, we map the 36 items to the 3 scales, which gives the inferred scores as shown in Table 1. As expected, the first sentence shows a high score for BOND, the second for TASK and GOAL, and the third for GOAL, demonstrating the model’s ability to differentiate and quantify the various aspects of the therapeutic working alliance.
@@ 2.5 Identification of Working Alliance Interaction Patterns
Specifically, we concatenate the 36-dimensional working alliance scores estimated from the current turn, as described above, with the unbiased sentence embedding of the turn. This combined feature vector is then fed into a sequence classifier, which we term Working Alliance Transformer (WAT) and Working Alliance LSTM (WA-LSTM) (see Supplementary Materials). By applying the WAT or the WA-LSTM to the psychotherapy transcripts, we can classify the clinical condition of the sequence based on the working alliance scores and the content of the dialogue turns. This classification model can be applied to the entirety of a session or a segment of the session; its accuracy provides a validation that these patterns can be identified, as we show in the Results section.
@@ 2.6 Psychotherapy Topic Modeling Framework
Using a language model to interpret text data offers several advantages for researchers and practitioners: (1) the model can provide an consistent analysis of the data that, while necessarily incomplete, is devoid of individual biases that analysts might inadvertently introduce; (2) by relying on the model’s interpretation, researchers can access a more neutral perspective, enhancing the objectivity of their findings; (3) language models can quickly process and analyze large volumes of text, identifying patterns, relationships, and insights that may be challenging for humans to detect manually, and eventually guide and supplement the daily practice of therapy.